Calculate the molarity, molality, and mole fraction of NH3 for a solution of 30.0 g NH3 in 70.0 g H2O. The density of the solution is 0.982 g/mL.

1
Expert's answer
2022-03-06T17:07:05-0500

Molarity = moles of solute / volume of solution = n / V

n(NH3) = m(NH3) / Mr(NH3) = 30.0 g / 17.02 g/mol = 1.76 mol

V(solution) = m(solution) / "\rho" (solution) = (30.0 g NH3 +O 70.0 g H2O) / 0.982 g/mL = 101.83 mL = 0.1018 L

M = 1.76 mol / 0.1018 L = 17.3 mol/L

Molality = moles of solute / kg of solvent

m = 1.76 mol NH3 / 0.070 kg H2O = 25.1 mol/kg

Mole fraction = moles of solute / moles of solution

n(H2O) = m(H2O) / Mr(H2O) = 70.0 g / 18.015 g/mol = 3.88565 mol

X(NH3) = 1.76 mol NH3 / (1.76 mol NH3 + 3.88565 mol H2O) = 0.3117